# Themis: Towards Flexible and Interpretable NLG Evaluation

### v1

@@ Abstract
The evaluation of natural language generation (NLG) tasks is a significant and longstanding research issue. With the recent emergence of powerful large language models (LLMs), some studies have turned to LLM-based automatic evaluation methods, which demonstrate great potential to become a new evaluation paradigm following traditional string-based and model-based metrics. However, despite the improved performance of existing methods, they still possess some deficiencies, such as dependency on references and limited evaluation flexibility. Therefore, in this paper, we meticulously construct a large-scale NLG evaluation corpus NLG-Eval with human and GPT-4 annotations to alleviate the lack of relevant data in this field. Furthermore, we propose Themis, an LLM dedicated to NLG evaluation, which has been trained with our designed multi-perspective consistency and rating-oriented preference alignment methods. Themis can conduct flexible and interpretable evaluations without references, and it exhibits superior evaluation performance on various NLG tasks, simultaneously generalizing well to unseen tasks and surpassing other evaluation models, including GPT-4.
Themis: Towards Flexible and Interpretable NLG Evaluation
@@ 1 Introduction
Automated evaluation is crucial for natural language generation tasks, as it measures the performance of related models and consequently promotes the development of NLG research. In the early years, traditional string-based evaluation metrics, such as BLEU (Papineni et al., 2002), were commonly used. Despite their convenience, surface-level matching cannot reliably evaluate texts as they are easily affected by perturbations (He et al., 2023), and previous work (Sulem et al., 2018a) has indicated their low correlation with human evaluations. With the development of pre-trained language models and related corpora, pre-trained model-based evaluation metrics such as BARTScore (Yuan et al., 2021) and COMET (Rei et al., 2020) have then been proposed. But their performance remains unsatisfactory compared with human evaluation, and they cannot conduct evaluations on customized aspects, like usefulness.
Figure 1: Themis is capable of evaluating various NLG tasks according to flexible evaluation aspects and criteria without references and providing corresponding analyses along with the evaluation ratings.
@@
To obtain high-quality training data, we extensively surveyed existing studies on NLG, collecting and filtering out 58 evaluation datasets with human annotations across 9 common NLG tasks. Given the importance of evaluation criteria, we meticulously proofread each dataset and manually supplemented the missing descriptions. In addition, GPT-4 is treated as an additional annotator to supplement and validate the collected data while providing evaluation analyses. Ultimately, we constructed a large-scale evaluation corpus, NLG-Eval, which contains about 0.5 million samples. These resources are released to alleviate the data scarcity in the field and facilitate relevant research.
Furthermore, we proposed a multi-perspective consistency verification method to select relatively more reliable data from the constructed NLG-Eval corpus. Empirical methods are also employed to increase the distribution balance and diversity of data, resulting in approximately 67K training samples. Moreover, we designed specific preference alignment, guiding the construction and utilization of preference data through evaluation ratings, to improve the evaluation capabilities of the fine-tuned model. Experimental results show that our Themis achieves better overall evaluation performance over previous evaluation models on common NLG tasks, including summarization, story generation, data-to-text, and so on. We also conducted more in-depth analyses, including generalization tests on unseen tasks like the instruction-following evaluation as well as aspect-targeted perturbation tests to verify the reliability of our model.
@@
  • We construct a large-scale NLG evaluation corpus, including about 0.5 million samples and 58 datasets across 9 NLG tasks, with detailed meta-information, aspect criteria, and evaluations.
@@
  • Extensive experiments demonstrate the superior evaluation performance of Themis in common NLG tasks, as well as good generalization and robustness. Our model and relevant resource are released in https://github.com/PKU-ONELab/Themis to facilitate related research.
@@ 2 NLG-Eval Corpus
Despite abundant data on NLG tasks, the corresponding high-quality evaluation data remains scarce and scattered due to the high cost of professional human annotations. Previous related methods either used a small amount of human annotations to train regression models, such as UniEval (Zhong et al., 2022) and X-Eval (Liu et al., 2023a), resulting in limited task coverage and a lack of evaluation analyses, or they entirely relied on LLMs to generate synthetic data, which raised reliability concerns (Hu et al., 2024). To address this challenge, we formally defined the NLG evaluation task and clarified the involved elements. Subsequently, we surveyed a large number of existing related studies and compiled 58 evaluation datasets with human annotations across 9 NLG tasks, totaling 0.5 million samples. They have undergone our meticulous proofreading, and the missing but critical content has been supplemented, such as evaluation criteria. Additionally, we utilize the competent GPT-4 for supplementary evaluations, including analyses and ratings. The corpus has also been equipped with much meta-information, aiming to promote the development of related research.
@@ 2.1 Definitions
NLG tasks typically require language models to generate the corresponding output $t$ based on the given input $i$ and task requirements. The input in specific tasks may consist of multiple components; for example, in controlled generation, it is composed of the generation requirements and the constraint labels. NLG evaluation involves evaluating $t$ in conjunction with $i$ through the evaluation model or method $E$, using evaluation criteria $c$ and references $r$ if available. When they are absent, the evaluation defaults to general and reference-free evaluation, respectively. Evaluation results come in various forms, primarily focusing on rating $s$ and additional information $a$ such as analyses. The formalization of the evaluation process is as follows, with square brackets indicating optional operations:
@@ 2.2 Data Collection
To improve the coverage and diversity of the corpus, we conducted a comprehensive survey of related work on NLG tasks, including summarization, dialogue response generation, data-to-text, and so on. We collected datasets with human evaluation ratings and performed meticulous manual checking and filtering, ultimately obtaining 58 high-quality datasets across 9 different NLG tasks, which contain approximately 0.5 million samples. However, we discovered that many datasets, despite having evaluation aspects, lacked specific criteria or definitions, which could lead to confusion and ambiguity in evaluations. Such concerns have also been pointed out by Zhou et al. (2022); Hu et al. (2024). So we analyzed the involved datasets, taking samples, human labels, and information from related papers into account, and manually supplemented the criteria.
Although human evaluation is often considered the most reliable, many datasets have sparse or inconsistent human annotations, leading to questions about their accuracy. Therefore, we utilized GPT-4 to cross-verify with human evaluations to improve reliability, as it has been shown to perform well on evaluation (Zheng et al., 2023). On the other hand, it can provide evaluation analyses beyond just ratings and increase interpretability, which is lacking in human evaluation. More implementation details can be found in Appendix A. Additionally, our subsequent training did not use references because many datasets lack references or only have automatically generated references that are unsatisfactory (Pu et al., 2023). And the references are also difficult to obtain in practical evaluations. The tasks included in our NLG-Eval corpus and their respective evaluation aspects are shown in Figure 2, with more statistics and detailed information presented in Appendix B.
@@ 3 Methodology
Existing studies have highlighted that data quality is more crucial than quantity in LLM fine-tuning (Li et al., 2023c; Zhou et al., 2023). Therefore, we adopted empirical methods to improve data diversity and distribution balance and designed a multi-perspective consistency verification approach to obtain about 67K high-quality samples from the constructed corpus. Then they are used for supervised fine-tuning based on open-source LLMs. To further enhance the evaluation capability and alignment of the model, we proposed a preference data construction and training method guided by evaluation ratings to improve the fine-tuned model.
@@ 3.1 Diversity and Balance
Similar to many studies on LLM training, we initially employed empirical approaches to sample data from the original corpus, ensuring adequate diversity and distribution balance, which is generally considered beneficial for model training and generalization. To be specific, we treated the NLG task, evaluation aspect, and rating triplets as identifiers of data categories to cover as comprehensive a distribution as possible. And we sampled 100 pieces of data from each category while also considering the semantic diversity of the texts involved. Given the insufficiency of data in some categories and the requirement to introduce diversity in evaluation analyses and aspect criteria, we conducted multiple samplings of analyses differently for each category to further balance the data distribution, simultaneously rephrasing the evaluation aspect criteria. In the end, we obtained a dataset of approximately 67K for supervised fine-tuning. The detailed sampling method for both samples and evaluation analyses will be introduced in the next subsection.
@@ 3.2 Multi-perspective Consistency Verification
During the sampling process, most categories actually had sufficient samples, and each sample also involved multiple annotated evaluations. To take full advantage of such sufficient data resources and enhance the reliability of the training data, we verified the consistency of each sample and their evaluations from three different perspectives, ultimately filtering out the potentially high-quality data.
@@
When dealing with complex tasks like mathematical problems and logical reasoning, LLMs often employ the chain of thought (CoT) (Wei et al., 2022) method. It involves generating intermediate steps before providing the final result, which generally leads to higher-quality responses. Furthermore, by performing multiple diverse samplings and selecting the most frequently occurring final result from the responses, the model’s performance can be further improved (Wang et al., 2023c). This is referred to as self-consistency, which can be calculated as the proportion of the most frequently occurring result to the total number of samplings. A higher self-consistency indicates greater certainty of the model itself and, thus, higher reliability of the corresponding responses. In our evaluation annotation, GPT-4 was required to generate an analysis before assigning a rating, similar to a CoT process, which allows using self-consistency for data filtering. So we first collected samples where at least half of the evaluation ratings are consistent, and such ratings are named consistent ratings. Our training data would be selected from these samples and their evaluations that include consistent ratings.
@@
The evaluation ratings serve as the most critical supervision signal in evaluation, making their accuracy paramount. Unlike previous related work that relied entirely on human labels or LLM-generated labels, we comprehensively considered the evaluations from both sources to conduct cross-validation. As for each sample, the corresponding consistent rating is regarded as its rating from GPT-4 and scaled to the same range as the human rating. We prioritized samples where the two evaluation ratings were close, which reflected the high consistency between evaluations from humans and GPT-4 and also indicated the potential strong reliability of the samples.
@@
Hu et al. (2024) indicated that although current LLMs possess exciting evaluation capabilities, they may encounter issues of confusing aspect criteria, affecting the accuracy of their analyses and ratings. To address this, we propose two specific criteria for inspecting the evaluations themselves: consistency between the evaluation analysis and rating, and consistency between the evaluation analysis and aspect. We designed additional instructions to prompt GPT-4 to re-evaluate the candidate evaluations, whose details could be found in Appendix A, and prioritized those performing well from both perspectives.
@@ 3.3 Preference Alignment
With the rise of InstructGPT (Ouyang et al., 2022), reinforcement learning from human feedback (RLHF), as one of the key technologies, has garnered wide attention and been applied in many subsequent LLMs with great success. So we have specifically modified DPO (Rafailov et al., 2023), a commonly-used implementation of RLHF, based on our evaluation scenarios, using evaluation ratings to guide the construction and training of preference data. The preference alignment was further conducted after the supervised fine-tuning of the model.
@@ 3.3.1 Preference Construction
Although the DPO method does not require training a separate reward model, it still necessitates preference pairs to convey preference information. Benefiting from multiple evaluations of each sample, we can directly regard evaluations that meet the consistent rating introduced in the Section 3.2 as the chosen response and those that do not as the rejected response to construct preference pairs. The quality gap between the two evaluations can be reflected by the difference in their evaluation ratings, as consistency ratings are deemed more reliable. And our preference data is constructed on the previously used training data without involving additional samples.
@@ 3.3.2 Rating-guided DPO
According to DPO, we can reparameterize the reward function $r$ using the policy as follows:
@@
where $\pi_{\theta}$ is the policy model, $\pi_{\mathrm{ref}}$ is the reference model, typically a model undergone supervised fine-tuning, $\beta$ is the parameter controlling the degree of deviation between the policy and the reference, and $Z(x)$ is the partition function. The Bradley-Terry (BT) (Bradley and Terry, 1952) model is employed to model preferences, yielding:
@@
$p^{*}(y_{1}\succ y_{2}|x)=\sigma(r(x,y_{1})-r(x,y_{2})-\alpha|R(y_{1})-R(y_{2})|)$
where $R(y)$ denotes the evaluation rating in $y$. Finally, the maximum likelihood of new preference model serves as our optimization objective:
@@
where $(x,y_{c},y_{r})$ contains the input, the chosen response, and the rejected response, respectively.

### v2

# Themis: A Reference-free NLG Evaluation Language Model with Flexibility and Interpretability

@@ Abstract
The evaluation of natural language generation (NLG) tasks is a significant and longstanding research area. With the recent emergence of powerful large language models (LLMs), some studies have turned to LLM-based automatic evaluation methods, which demonstrate great potential to become a new evaluation paradigm following traditional string-based and model-based metrics. However, despite the improved performance of existing methods, they still possess some deficiencies, such as dependency on references and limited evaluation flexibility. Therefore, in this paper, we meticulously construct a large-scale NLG evaluation corpus NLG-Eval with annotations from both human and GPT-4 to alleviate the lack of relevant data in this field. Furthermore, we propose Themis, an LLM dedicated to NLG evaluation, which has been trained with our designed multi-perspective consistency verification and rating-oriented preference alignment methods. Themis can conduct flexible and interpretable evaluations without references, and it exhibits superior evaluation performance on various NLG tasks, simultaneously generalizing well to unseen tasks and surpassing other evaluation models, including GPT-4.
Themis: A Reference-free NLG Evaluation Language Model with Flexibility and Interpretability
@@ 1 Introduction
Automated evaluation is crucial for natural language generation tasks, as it measures the performance of related models and consequently promotes the development of NLG research. In the early years, traditional string-based evaluation metrics, such as BLEU (Papineni et al., 2002), were commonly used. Despite their convenience, surface-level matching cannot reliably evaluate texts as they are easily affected by perturbations (He et al., 2023), and previous work (Sulem et al., 2018a) has indicated their low correlation with human evaluations. With the development of pre-trained language models and related corpora, pre-trained model-based evaluation metrics such as BARTScore (Yuan et al., 2021) and COMET (Rei et al., 2020) have then been proposed. But their performance remains unsatisfactory compared with human evaluation, and they cannot conduct evaluations on customized aspects, like coherence.
Figure 1: Themis is capable of evaluating various NLG tasks based on flexible evaluation aspects and criteria without references and providing corresponding analyses along with the evaluation ratings.
@@
To obtain high-quality training data for NLG evaluation, we conducted a comprehensive survey of relevant studies and collected corresponding resources, finally selecting 58 evaluation datasets with human annotations across 9 common NLG tasks. Given the importance of evaluation criteria, we meticulously proofread each dataset and manually supplemented the missing descriptions. In addition, GPT-4 is treated as an additional annotator to supplement and validate the collected data while providing evaluation analyses. Ultimately, we constructed a large-scale evaluation corpus, NLG-Eval, which contains about 0.5 million samples with meta information. It aims to alleviate the issue of scattered and scarce data in the area of NLG evaluation and facilitate relevant research.
Furthermore, we proposed a multi-perspective consistency verification method to select relatively more reliable data from the constructed NLG-Eval corpus. Empirical methods are also employed to ensure sufficient diversity and balanced distribution of the data as much as possible, resulting in approximately 67K training samples. Moreover, we designed specific preference alignment, guiding the construction and utilization of preference data through evaluation ratings, to improve the evaluation capabilities of the fine-tuned model. Experimental results show that our Themis achieves better overall evaluation performance over previous evaluation models on common NLG tasks, including summarization, story generation, and so on. We also conducted more in-depth analyses, including generalization tests on unseen tasks like the instruction-following evaluation as well as aspect-targeted perturbation tests to verify the reliability.
@@
  • We construct a large-scale NLG evaluation corpus, including about 0.5 million samples and 58 datasets across 9 NLG tasks, with detailed meta information, aspect criteria, and evaluations from both humans and GPT-4.
@@
  • Extensive experiments demonstrate the superior evaluation performance of Themis on common NLG tasks, as well as good generalization and robustness. Our model and relevant resource have been released in https://github.com/PKU-ONELab/Themis to facilitate related research.
@@ 2 NLG-Eval Corpus
Despite abundant data on NLG tasks, the corresponding high-quality evaluation data remains scarce and scattered due to the high cost of professional human annotations. Previous related methods either used a small amount of human annotations to train regression models, such as UniEval (Zhong et al., 2022) and X-Eval (Liu et al., 2023a), resulting in limited task coverage and a lack of evaluation analyses, or they entirely relied on LLMs to generate synthetic data, which raised reliability concerns (Hu et al., 2024). To address this challenge, we formally defined the NLG evaluation task and clarified the involved elements. Subsequently, we surveyed a large number of existing related studies and compiled 58 evaluation datasets with human annotations across 9 NLG tasks, totaling 0.5 million samples. They have undergone our meticulous proofreading, and the missing but critical content has been supplemented, such as evaluation criteria. Additionally, we utilize the competent GPT-4 for supplementary evaluations, including analyses and ratings. The corpus has also been equipped with meta information, aiming to promote the development of related research.
@@ 2.1 Definitions
NLG tasks typically require language models to generate the corresponding output $t$ based on the given input $i$ and task requirements. The input in specific tasks may consist of multiple components; for example, in controlled generation, it is composed of the generation requirements and the constraint labels. NLG evaluation involves evaluating $t$ in conjunction with $i$ through the evaluation model or method $E$, using evaluation criteria $c$ and references $r$ if available. When they are absent, the evaluation defaults to general and reference-free evaluation, respectively. Evaluation results come in various forms, primarily focusing on rating $s$ and additional information $a$ such as analyses. The formalization of the evaluation process is as follows, with square brackets indicating optional elements:
@@ 2.2 Data Collection
To improve the coverage and diversity of the corpus, we conducted a comprehensive survey of related work on NLG tasks, including summarization, dialogue response generation, data-to-text, paraphrase, and so on. We collected datasets with human evaluation ratings and performed meticulous manual checking and filtering, ultimately obtaining 58 high-quality datasets across 9 different NLG tasks, which contain about 0.5 million samples. However, we discovered that many datasets, despite including evaluation aspects, lacked specific criteria or definitions, which could lead to confusion and ambiguity in evaluations. Such concerns have also been pointed out by Zhou et al. (2022); Hu et al. (2024). So we analyzed all these datasets, taking samples, human labels, and information from related papers into account, and manually supplementing the criteria. The different tasks included in our NLG-Eval corpus and their respective involved evaluation aspects are gathered and shown in Figure 2, with semantically similar aspects being merged. And more detailed statistics and information are presented in Appendix A.
2.3 GPT-4 Annotations
Although human evaluation is often considered the most reliable, many datasets have sparse or inconsistent human annotations, leading to questions about their accuracy. Therefore, we utilized the superior language capabilities of GPT-4 to evaluate each sample in the corpus, as it has been shown to perform well on evaluation (Zheng et al., 2023). GPT-4 can serve as an additional annotator to cross-verify with human evaluations, thereby improving reliability. On the other hand, it can provide evaluation analyses beyond just ratings and increase interpretability, which is lacking in human evaluation. Following conclusions and suggestions from Chiang and Lee (2023b); Bsharat et al. (2023), we carefully designed instructions, ensuring that a concise and accurate analysis is provided before the rating. Each sample is evaluated with a temperature setting of 1, and 10 diverse evaluation results are obtained through multiple samplings, with more details included in Appendix B. Additionally, our annotation and subsequent training did not use references because many datasets lack references or only have automatically constructed references that are unsatisfactory (Pu et al., 2023). And the references are actually difficult to obtain in practice.
@@ 3 Methodology
Existing studies have highlighted that data quality is more crucial than quantity in LLM fine-tuning (Li et al., 2023c; Zhou et al., 2023). Therefore, we adopted empirical methods to improve the diversity and balance of data distribution and designed a multi-perspective consistency verification approach to obtain high-quality samples from the constructed corpus. Then they would be used for supervised fine-tuning based on open-source LLMs. To further enhance the evaluation capability and alignment of the model, we proposed a preference data construction and training method guided by evaluation ratings to improve the fine-tuned model.
@@ 3.1 Diversity and Balance
Similar to many studies on LLM training, we initially employed empirical approaches to sample data from the original corpus, ensuring adequate diversity and balanced distribution, which was generally considered beneficial for model training and generalization. To be specific, we treated the NLG task, evaluation aspect, and rating triplets as identifiers of data categories to cover as comprehensive a distribution as possible. And we sampled 100 pieces of data from each category while also considering the semantic diversity of the input texts involved. Given the insufficiency of data in certain categories and the requirement to introduce diversity in evaluation analyses and aspect criteria, we conducted multiple samplings of analyses differently for each category to further balance the data distribution, simultaneously rephrasing the evaluation aspect criteria. In the end, we obtained a dataset of approximately 67K for supervised fine-tuning. The detailed sampling method for both samples and evaluations will be introduced in the next subsection.
@@ 3.2 Multi-perspective Consistency Verification
During the sampling process, most categories actually include abundant samples, and each sample also involves multiple evaluations annotated by GPT-4. To take full advantage of such sufficient data resources and enhance the reliability of the training data, we verified the consistency of each sample and their evaluations from three different perspectives, ultimately obtaining the potentially high-quality data through screening.
@@
When dealing with complex tasks like mathematical problems and logical reasoning, the chain of thought (CoT) (Wei et al., 2022) method is always employed on LLMs to improve their performance. It involves generating intermediate steps before providing the final result, which generally leads to higher-quality responses. Furthermore, by performing multiple diverse samplings and selecting the most frequently occurring final result from the responses, the output can be more stable and accurate (Wang et al., 2023c). This process is referred to as self-consistency, which can also be regarded as a metric and calculated as the proportion of the most frequently occurring result to the total number of samplings. A higher self-consistency indicates greater certainty of the model itself and, thus, higher reliability of the responses to the corresponding samples.
In our evaluation annotation, GPT-4 was required to generate an analysis before assigning a rating, similar to a CoT process, which allows using self-consistency for data filtering. The formal definitions of the most frequently occurring rating $\hat{r}$ (called consistent rating) and self-consistency are as follows, where $n$ denotes the number of evaluations for a sample, and $r_{i}$ represents the rating of the $i$-th evaluation:
$\hat{r}=\arg\max_{r}$ $\sum_{i=1}^{n}\mathbbm{1}\left(r_{i}=r\right)$
[…]
And we prioritized samples with high self-consistency and only retained their evaluations that include the consistent rating $\hat{r}$ to collect both potentially reliable samples and evaluations for the following process of verification.
@@
The evaluation ratings serve as the most critical supervision signal in evaluation, making their accuracy paramount. Unlike previous related work that relied entirely on human labels or LLM-generated labels, we comprehensively considered the evaluations from both sources to conduct cross-validation. As for each sample, the corresponding consistent rating $\hat{r}$ is regarded as its rating from GPT-4 and scaled to the same range as the human rating. We prioritized samples where the two evaluation ratings were close, which reflected the high consistency between evaluations from humans and GPT-4 and also indicated the potential strong reliability of the samples.
@@
Hu et al. (2024) indicated that although current LLMs like GPT-4 possess exciting evaluation capabilities, they may encounter issues of confusing aspect criteria, affecting the accuracy of their analyses and ratings. To address this, we propose two specific criteria for inspecting the evaluations themselves: consistency between the evaluation analysis and rating, and consistency between the evaluation analysis and aspect. We designed additional instructions to prompt GPT-4 to re-evaluate the candidate evaluations, whose details could be found in Appendix B, and prioritized those being assessed as of good quality from both perspectives.
@@ 3.3 Preference Alignment
With the rapid rise of InstructGPT (Ouyang et al., 2022), reinforcement learning from human feedback (RLHF), as one of the key technologies, has garnered wide attention and been applied in many subsequent LLMs with great success. So we have specifically modified DPO (Rafailov et al., 2023), a commonly-used implementation of RLHF, based on our NLG evaluation scenarios, using evaluation ratings to guide the construction and training of preference data. The preference alignment was further conducted after the supervised fine-tuning of the model.
@@ 3.3.1 Preference Construction
Although the DPO method does not require training a separate reward model, it still necessitates preference pairs to convey preference information. Benefiting from multiple evaluations of each sample, we can directly regard the evaluation that matches the consistent rating $\hat{r}$ introduced in Section 3.2 as the chosen response and that does not as the rejected response to construct preference pairs. The quality gap between the chosen and rejected evaluations can be reflected by the difference in their evaluation ratings, as the closer the rating to $\hat{r}$, the more reliable the evaluation can be deemed. And our preference data is constructed based on the previously obtained training data without involving additional samples.
@@ 3.3.2 Rating-guided DPO
During vanilla DPO, we reparameterize the reward function $r$ using the policy as follows:
@@
where $\pi_{\theta}$ is the policy model, $\pi_{\mathrm{ref}}$ is the reference model, typically a model undergone supervised fine-tuning, $\beta$ is the parameter controlling the degree of deviation between them, and $Z(x)$ is the partition function. Moreover, the Bradley-Terry (BT) (Bradley and Terry, 1952) model is employed to model preferences, yielding:
@@
$p^{*}(y_{1}\succ y_{2}|x)=\sigma($ $r(x,y_{1})-r(x,y_{2})$
$-\alpha|R(y_{1})-R(y_{2})|)$
where $R(y)$ denotes the evaluation rating included in response $y$. Finally, the maximum likelihood of the new preference model serves as the optimization objective for our rating-guided DPO:
@@
where $(x,y_{c},y_{r})$ contains the input content, the chosen evaluation, and the rejected evaluation, respectively.